oslo company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units): sales $85,000, variable expenses 59,500, contribution margin 25,500, fixed expenses 20,400, net operating income $5,100. what is the degree of operating leverage? ( round answer to 2 decimal places.) using the degree of operating leverage, what is the estimated percent increase in net operating income of a 5% increase in sales? (round intermediate calculations and final answer to 2 decimal places.) assume that the amounts of the company's total variable expensed and total fixed expenses were reversed. In other words, assume that the total variable expenses are $20,400 and the total fixed expensed are $59,500. under this scenario and assuming that total sales remain the same, what is the degree of operating leverage? (round answer to 2 decimal places.) Assume that the amounts of the company's total variable expenses and total fixed expenses were reversed. in other words, assume that the total variable expenses are $20,400 and the total fixed expenses are $59,500. Given this scenario and assuming that total sales remain the same. using the degree of calculated operating leverage, what is the estimated percent increase in net operating income of a 5% increase in sales? (round your intermediate calculation and final answer to 2 decimal places.)
Answer #1

1. Degree of operating leverage = Contribution Margin / Net operating income = 25500/ 5100 = 5

2. Estimated percent increase in net operating income of a 5% increase in sales = 5% *5 = 25%

3.   Degree of operating leverage = Contribution Margin / Net operating income = 64600/ 5100 = 12.67

Sales 85,000
Less: variable expenses 20400
Contribution margin 64,600
Fixed expenses 59500
Net operating income 5,100

4. Estimated percent increase in net operating income of a 5% increase in sales = 12.67 *5% = 63.33%